40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 322 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  898   Fri Aug 29 11:05:11 2008 josephbSummaryComputersc1asc was down this morning
I had to manually reboot c1asc this morning, as for some unknown reason its status was red, and the fiber lights on the board were status:red, sig det:amber, own data: nothing. Shut the crate down, turned it back on, heard a beep, then followed wiki reboot instructions. Seems to be working now.
  897   Fri Aug 29 11:01:49 2008 josephbConfigurationComputersAttempt to change a channel gain in ICS-110B
As noted earlier by Rana, I was playing around with the /cvs/cds/caltech/chans/daq/C1IOOF.ini file with help from Rob. I had made a backup before hand and saved it as C1IOOF.ini.Aug-28-2008. (I have since been informed that C1IOOF.ini.082808 would have been prefered as a name).

We had been trying to up the gain in the C1: PSL-ISS_INMONPD_F in order to do a very low power PMC sweep, in an attempt to get clean modes for fitting. Initially we pressed the reconfig button on the C0DAQ_DETAIL screen, but all that seemed to do was change the Config File CRC. We proceeded to reboot fb40m remotely. However, any change to the ini file (even an extra space at the end of the file) caused a 0x2000 status for C1IOVME16k on the C0DAQ_DETAIL screen. At the time I presumed it was comparing the CRC of the ini-file to something else.

Digging around on in Alex's webspace at http://www.ligo.caltech.edu/~aivanov/ , I found the NDS Access page, which indicated that 0x2000 was a conflict between the front-end and frame builder .ini files.

"There is also status bit 0x2000 which gets added when the DCU configuration is different in front-end and frame builder. That is you can change and .ini file an then reload DAQ configuration with Epics button, which reconfigures the front-end, but leaves frame builders with invalid old configuration. They will detect this change and set the status to 0x2000 to indicate this condition. You will have to restart frame builders to pick up new .ini file and set status back to zero for the affected DCU."

It was when I was going to try reseting the c1iovme via the C0DAQ_RFMNETOWRK medm screen that we realized the EPICS controls were not responding properly. The .ini file was returned to its original form, and mass reboots commenced.
  896   Fri Aug 29 10:20:32 2008 YoichiConfigurationPSLbeam block distorted

Quote:
There was a beam block after the Mach Zender. Who or what put this there?

The going to the MC now looks distorted as if someone has left something funny in the beam or maybe the new PMC has started to degrade??

Use the ELOG people...its good for you.


I put the block. I was frequently reaching to the FSS box to change the test point probes. I put the block to protect my hands/clothes from being burnt accidentally.
  895   Fri Aug 29 02:40:43 2008 rana,jenneUpdatePSLPMC Servo Board

Quote:
Board is back in. PMC is locked.


This entry has details about the low pass filter after the PMC mixer. This filter has a few purposes:

1] Remove the beat signal (at 2*f_mod) between the PD RF signal at f_mod and the LO signal at f_mod.
2] Remove the beat signal (at f_mod) between the PD RF signal at 2*f_mod (which comes from the
beating of the upper and lower RF sidebands) and the LO signal at f_mod.
3] Remove other RF signals from non-ideal behavior of the LO drive signal and distortion in the RF PD pre-amp.


So its important to have a very good rejection at 35 MHz and higher. I used the Hartmut LC network design which is
installed on H1, H2, & L1. Since there is a high gain in the audio amps right after the mixer we have to get rid of
the RF or else we'll get slew rate limited or otherwise rectified downconversion of the RF signal into our audio band.

Of course, what everyone immediately realizes from the above 3 points, is that this filter can't protect the PMC
noise performance from homodyne mixing (e.g. 2*f_mod in the LO and 2*f_mod in the RF PD). To get around that, we're
ordering some filters from Mini-Circuits to remove the 2f from those signals by ~30 dB. As long as we install
the same filters on the RF and LO legs, there should be no significant phase shift in the demodulated signal.

The attached 2 page PDF shows the calculated before and after TFs of this filter. The 2 attached .m files
calculate the TF's and have ascii art which shows how the filter works.

Here's a comparison of the attenuation (in dB) of 2 candidate Mini-circuits filters:

f(MHz)SLP-30SLP-50
31 0.5 0.4
35 1.3 0.4
38 6.1 0.4
40 10.8 0.42
61 46.3 14.8
71 60 29
91 76.9 48
10780 60

We don't have tabulated data at the same frequencies for both filters so I just made up some of the points by eye-balling the
plots from the catalog - but you get the idea: we can get away with using the SLP-30 at 35 MHz since it only attenuates the
signals by ~1.5 dB. So if someone can find 4 of these then Steve doesn't have to order any from Mini-Circuits.
Attachment 1: pmclp-07.pdf
pmclp-07.pdf pmclp-07.pdf
Attachment 2: pmclp_40m_080824.m
% PMCLP is a TF of the IF filter after the PMC mixer
%       

% Mixer_Voltage -- Rs -- L1 --- L2 ---------Vout
%                            |      |   |
%                           C1     C2   Rl                   
%                            |      |   |
%                           GND    GND  GND
%

... 58 more lines ...
Attachment 3: pmclp.m
% PMCLP is a TF of the IF filter after the PMC mixer
%       

% Mixer_Voltage -- Rs -- L1 --- L2 ---------Vout
%                            |      |   |
%                           C1     C2   Rl                   
%                            |      |   |
%                           GND    GND  GND
%

... 57 more lines ...
  894   Thu Aug 28 19:02:25 2008 rana, josephb, robSummaryComputersbig boot
This afternoon Joe did something with an .ini file (look for his detailed elog entry) and the computers went bad.
RFM network screen not active - filter modules not working.

We went around and booted every machine as has been done before. The correct order for a memory corruption
fixing big boot is the following:

    [1] RESET the RFM switches near the FB racks.
    [2] Power cycle c1dcuepics.
    [3] Power cycle all other crates with real time CPUs:
    c1iscey, daqctrl, daqawg, c1susvme1, c1susvme2, c1sosvme, c1iovme, c1lsc, c1asc, & c1iscex
    [4] Start up all FEs as described in Wiki.
    [5] Burt restore everyone (losepics, iscepics, assepics, omcepics?)
  893   Thu Aug 28 18:56:14 2008 ranaConfigurationPSLbeam block distorted
There was a beam block after the Mach Zender. Who or what put this there?

The going to the MC now looks distorted as if someone has left something funny in the beam or maybe the new PMC has started to degrade??

Use the ELOG people...its good for you.
  892   Wed Aug 27 13:55:43 2008 rana,jenneUpdatePSLPMC Servo Board
Board is back in. PMC is locked.

Nominal gain is now 15 dB with brick. We need to do more studies:

  • Find out why there is still 35 MHz signal at the error point. Order some low pass filters to cut off above 35 MHz.
  • Explore brick + no-brick loop shapes and error spectra.
  • Measure and set the OLG.

We've left the copper-wrapped lead brick installed to let it slowly conform to the glass better.
  891   Wed Aug 27 12:09:10 2008 EricSummaryCamerasWeekly Summary
I added a configuration file parser to the Snap code. This allows all command line parameters (like exposure time, etc.) to be saved in a file and loaded automatically. It also provides a method of loading parameters to transform a point from its location on the image to its location in actual space (loading these parameters on the command line would substantially clutter it). The code is now fully set-up to test servo-ing one of the mirrors again, and I will test this as soon as the PMC board stops being broken and I can lock the X-arm.

I also took an image of the OSEMs on ETMX in order to apply the rotation transform code in order to determine the parameters to pass to Snap. The results were alpha = 2.9505, beta = 0.0800, gamma = -2.4282, c = 0.4790. These results are reasonable but far from perfect. One of the biggest causes of error was in locating the OSEMs: it is difficult to determine where in the spot of light the OSEM actually is, and in one case, the center was hidden behind another piece of equipment. Nevertheless, the parameters are good enough to use in a test of the ability to servo, though it would probably be worth trying to improve them before using them for other purposes. The original and rotated images are attached.

I've begun working on calculations to figure out how much power loss can occur due to a given cavity misalignment or change in a mirror's radius of curvature from heating. The goal is to determine how well a camera can indirectly detect these power losses, since a misalignment produces a change in beam position and a change in radius of curvature produces a change in beam waist, both of which can be measured by the camera.

Joe and I hunted down the requisite equipment to amplify the photodiode at the output of the PMC, allowing us to turn the laser power down even more during a scan of the PMC, hopefully avoiding thermal effects. This measurement can be done once the PMC works again.
Attachment 1: originalETMX.png
originalETMX.png
Attachment 2: rotatedETMX.png
rotatedETMX.png
  890   Wed Aug 27 10:55:35 2008 YoichiHowToComputersAnnoying behavior of the touch pads of the lab. laptops is fixed
I was sick of the stupid touch pad behavior of the lab. laptops, i.e. firefox goes back and forth in the history when the cursor is moved.
It was caused by firefox mis-interpreting the horizontal scroll signal as back/forward command.
I stopped it by going to about:config in firefox and set mousewheel.horizscroll.withnokey.action to 0 and
mousewheel.horizscroll.withnokey.sysnumlines to true.
  889   Tue Aug 26 19:07:37 2008 YoichiHowToComputersReading data from Agilent 4395A analyzer through GPIB from *Linux* machine
I succeeded in reading data from Agilent 4395A analyzer, who's floppy is crappy, through GPIB from a Linux machine using
agilent 82357B USB-GPIB interface.
I installed the linux GPIB driver to one of the lab. laptops (the silver DELL one currently sitting on the 4395A analyzer).
I wrote an initialization script for the USB-GPIB interface and a small python script for reading data from the analyzer.

[Usage]

1. Connect the USB-GPIB interface to the laptop and the analyzer.
2. Run /usr/local/bin/initGPIB command (it takes about 10sec to complete).
3. Run /usr/local/bin/getgpibdata.py > data.txt to save data from the analyzer to a text file.

The data format is explained in the comments of getgpibdata.py
This method is way faster than the unreliable floppy. The data is transfered in a few sec.

I'm now writing a wiki page on this
http://lhocds.ligo-wa.caltech.edu:8000/40m/GPIB

I will install the same thing into the other DELL laptop soon.
Let me know if you have trouble with this.
  888   Tue Aug 26 18:19:16 2008 ranaOmnistructureElectronicsResistor Noise at the 40m
As Stefan points out in his recent ISS ilog entries at LLO, Daniel Sigg recently wrote a
recommendation memo on resistor and capacitor
choices: T070016.

While working on the PMC I have had to use leaded resistors and wondered about the noise. As it turns
out we have the RN series of 1/4 W resistors from Stackpole Electronics. The RN series are
metal film resistors (datasheet attached); metal film is what Sigg recommends for lowest flicker
noise.

So we are OK for using the Stackpole 1/4 W leaded resistors in low noise circuits.
Attachment 1: SEI-RN_RNM.pdf
SEI-RN_RNM.pdf SEI-RN_RNM.pdf
  887   Tue Aug 26 15:06:16 2008 steveConfigurationVACrga scan
Pumpdown 66 PRM-maglev vac normal -day 11
short form: pd66PRM-m-d11
Attachment 1: RGA-0808260125.png
RGA-0808260125.png
  886   Tue Aug 26 12:00:45 2008 JenneSummaryPEMTransfer function of Ranger seismometer
This finishes up the calibration that Rana started in elog # 881.

The calibration of the Ranger seismometer should also include:
2 zeros at 0 Hz
2 poles at 1.02 Hz

This comes from finding the transfer function between the mass's motion and the motion of the ground.
    ..
m * x  = (x_G - x) * k  + d(x_G - x) * b
                          dt

where
  • m = mass
  • x = displacement of the mass
  • x_G = displacement of the ground
  • k = spring constant
  • b = damping constant

This gives
x               w0^2  +  i*w*w0/Q
----    =    -----------------------
x_G           w0^2 + i*w*w0/Q - w^2

where
  • w0 = sqrt(k/m) = natural frequency of spring + mass
  • w = frequency of ground motion
  • Q = q-factor of spring + mass system = 1/2 for critically damped system

The readout of the system is proportional to
d  (x - x_G)          (    w0^2  +  i*w*w0/Q          )    .                    w^2               .
dt                 =  (  -----------------------  - 1 ) * x_G   =      ----------------------- * x_G
                      (   w0^2 + i*w*w0/Q - w^2       )                w0^2 + i*w*w0/Q - w^2
Since we read out the signal that is proportional to velocity, this is precisely the transfer function we're looking for. With w0 = 1.02 Hz and Q = 1/2 for the critically damped system, we have 2 zeros at 0 and 2 poles at 1.02.
  885   Tue Aug 26 09:58:59 2008 steveOmnistructureCOCETMX is #03
This is the picture of ETMX from the upper south west viewport
Attachment 1: etmx-d40posg_012.jpg
etmx-d40posg_012.jpg
Attachment 2: etmx03.jpg
etmx03.jpg
  884   Tue Aug 26 09:04:59 2008 ranaConfigurationPSLPMC Servo Board: Out for Repairs
I've started modifying our PMC board to bring it up to the 21st century - leave the screen alone or else you might zap something.
  883   Mon Aug 25 21:15:23 2008 ranaConfigurationLSCaux NPRO off
Looks like no one has used the Lightwave NPRO on the AS table after Koji left, so I turned it off so that it can rest until Alberto does the X-arm measurements.
  882   Mon Aug 25 17:45:34 2008 rana, josephb, robHowToPEMAccelerometer range
Joe shows us by jumping up ~15" in the control rom that the accelerometers are set with not enough gain.

Since this is taken around 5:30 in the evening, so we can take the nearby time series to represent what a
high noise level is. I recommend we up the gain using the ICS-110B .ini file.
Attachment 1: Screenshot-4.png
Screenshot-4.png
  881   Mon Aug 25 15:50:18 2008 ranaSummaryPEMRanger SS-1
The manual for the Ranger SS-1 seismometer can be found on line here:
ftp://ftp.kmi.com/pub/software_manuals/300190/300190nc.pdf

and now in our 40m PEM Wiki page:
Ranger_SS-1

To calibrate it, we use the formula from the manual:
                 R_x
G_L = G_0 * ------------   =  149 +/- 3 V/(m/s)
             R_x  +  R_c

where
G_0 = 340 V/(m/s)    (generator constant)
R_x = 4300 Ohms      (external damping resistor in Pomona box)
R_c = 5500 Ohms      (internal coil resistance)

Then we have a gain of 200 in the SR560 so that gets us to ~30000 V/(m/s).

And then there's a DAQ conversion factor of the usual 2^16 cts / 4 V.

so the calibration constant is

G = 488 counts / (micron/sec)

in the ~1-50 Hz band
  880   Mon Aug 25 14:42:09 2008 EricConfigurationCamerasETMX Digital Camera
I changed the lens on the camera looking at the ETMX to a 16mm, 1:1.4 zoom lens. This is in preparation to measure a couple parameters that depend on the camera's position and angle, so please avoid repositioning it for a couple of days.
  879   Mon Aug 25 14:18:36 2008 JenneUpdatePSLPMC servo board is fixed
The PMC servo board is back in place, all fixed up with a shiny new resistor. The PMC locks, and the MC locks (I'm not saying anything either way about how long the MC will stay locked, but it is locked for now). The resistor is connected to the connector using a short piece of wire, so this problem won't happen again, at least with this connector on this board.
  878   Mon Aug 25 12:13:49 2008 JenneUpdatePSLBroken PMC Servo Board
I broke the PMC servo board (on accident).

I was trying to measure the resistance of the extra resistor that someone put between the board and the HV OUT connector, since this is part of an RC filter (where C is the capacitance of the PZT on the PMC) that I need to know the values of as part of my mission to make a 14.6kHz notch for the PMC body mode. The resistance is 63.6k. I had to pull the board to get in to measure this resistance.

This resistor between the board and the center pin of the panel-mount HV OUT connector made a rigid connection between the board and the panel. When I was putting the board back in, I must have strained this connection enough that it broke. We don't have any of the same kind of resistor here at the 40m, so I'm waiting until after lunch to go to Wilson house and see if they've got any. The IFO is down until I get this sorted out.
  877   Mon Aug 25 11:43:55 2008 YoichiFrogsIOOMC REFL PD cable had been disconnected through out the weekend
Most of my morning was wasted by the MC REFL PD cable, which was disconnected on the generic LSC PD interface board.
I know who did this. *ME*. When I pulled out the MC board, which is sitting next to the PD interface, on Friday, I must have
disconnected the PD cable accidentally. The connector of the PD cable (D-Sub) does not have screws to tighten and easily comes off.
I wrote this entry to warn other people of this potential problem.
  876   Mon Aug 25 10:51:06 2008 steveUpdatePSLpsl headtemp is coming down
The laser water cooler was overflowing this morning.
I removed 500 cc water from the chiller.

The 4 days plot shows clearly:
that the capacity of the chiller is depending on the water level.
Overflowing water is a heat load for the chiller, so laser head temp goes up.
Attachment 1: ht4dw.jpg
ht4dw.jpg
  875   Mon Aug 25 10:23:53 2008 steveHowToGeneralcable killer
Rack 1Y7 double violation:

BNC cables left to be jammed by door

and see destroyed BNCs

RED fibers should be rerouted.
I placed protective obstacle in position
so the door can not be closed.

Please do not do this!

DNA analysis is in progress on your finger prints.
Attachment 1: cablkill.png
cablkill.png
Attachment 2: cablkll2.png
cablkll2.png
  874   Mon Aug 25 10:07:35 2008 JenneUpdatePSLNumbers for the PMC servo board (Re: entry # 873)
Jenne, Rana

These are the numbers that go along with Rana's entry #873:

The existing notch in the PMC servo is at 31.41kHz.

The power spectrum of the PMC has a peak at 14.683kHz when it is just sitting on the PSL table (no extra mass). When we put a pile of steel and aluminum (~20lbs) on top of the PMC, the body resonance moves to 14.622kHz, but is decreased by about 40 dB!

Rana has ordered a lead brick + foil that should arrive sometime this week. To complete the mechanical part of this installation, we need to extend the earthquake mounts around the PMC so that the lead brick can't fall off of the PMC onto the rest of the table.
  873   Sat Aug 23 09:39:51 2008 rana, jenneUpdatePSLPMC Survey
Jenne, Rana

We scoped out the PMC situation yesterday.

Summary: Not broke. UGF ~ 500 Hz. Needs some electronics work (notches, boosts, LPFs)

Ever since we swapped out the PMC because of the broken PZT of the previous one, the UGF has been
limited to a low value. This is because the notches no longer match the mechanical resonant
frequencies of the body. The old one had a resonance at 31.3 kHz which we were notching using
the LC notch on the board as well as a dangling Pomona box in the HV line to the PZT. The one
has a resonance at ~14.5 kHz which we don't yet have a notch for. Jenne has all the real numbers and
will update this entry with them.

Todo:
  • Implement the 4th order Grote low pass after the mixer.
  • Replace the AD797 with an OP27.
  • Change servo filter to have a boost (need DC gain)
  • Make a 14.5 kHz notch for the bode mode.
  • Put a 20 lb. gold-foil wrapped lead brick on the PMC.

Here's the link about the modified PMC board which we installed at LHO:
LHO PMC elog 2006
  872   Fri Aug 22 17:03:41 2008 YoichiUpdateIOOMC open loop TF
I measured the open loop TF of the overall MC loop using the sum-amp A of the MC board.
I used the Agilent 4395A network analyzer and saved the data into a floppy disk. However, the data was corrupted when
I read it with my computer. I had the same problem before. The floppy is not reliable. Anyway, I have to re-measure the TF.
From what I remember, the UGF was around 25kHz and the phase margin was less than 15deg.
Above this frequency, the open loop gain was almost flat and had a small bump around 100kHz.
This bump has a gain margin of less than 4dB (the phase is more than 180deg delayed here).
So the MC is marginally stable and either decreasing or increasing the gain will make it unstable easily.
Probably, the broken FSS is responsible for this. We have to fix it.

During the measurement, I also found that the input connectors (IN1 and IN2) of the MC board are freaky.
These are TNC connectors directory mounted on the board. Gently touching the cables hooked up to those connectors
caused a large offset change in the output.
When Rana pulled the board out and pushed it in firmly, the strange behavior went away. Probably, the board was
not correctly inserted into the backplane.
This could have been the reason for the MC unlocks.
  871   Fri Aug 22 16:06:29 2008 steveUpdatePEMparticle counter replaced, flowbenches & HEPAs checked
MetOne #2 counter was swapped in (on the top of IOC, facing SW direction, at ~75 deg upwards)
with channel one size 0.3 micron and channel two size 1.0 micron.
Sampling time was reduced from 60s to 6 sec at 0.1 cf/min at 25 min rate.

This means that displayed number needs to be multiplied by x100 to get particles/cf/min


HEPA filters and flow benches were checked:

PSL enclosure closed, HEPA speed at 60% 0-0 particles on optical table NW corner

AP covered optical table 1,000 particles of 0.3micron and 10 of 1.0 micron at NE corner

Flow bench at SE 0-0 particle (p)

on the top of SP cover at SE corner 60,000 p of 0.3 micron and 530 p of 1.0 micron

Mobile HEPAs 10cm from output screen in the center 800 p of 0.3 micron and 0 p of 1 micron
These filters will be replaced.

Clean assembly room:
both flow benches 0-0 p for 0.3-1.0 micron
east side bench 520 p of 0.3 micron and 210 p of 1.0 micron

Large hood in baking room with fan on 1.7 million p of 0.3 micron
and 16,000 p of 1.0 micron

Pasadena air just outside of main entrance:
3 million p of 0.3 micron and 30,000 p of 1.0 micron

My desk 743,000 p of 0.3micron, 63,000 p of 0.5 micron and 5,500 p of 1.0 micron cf/min

NOTE: existing COCHECKLIST.adl PEM displays needs to be corrected so it shows the 10 fold increase
and change particle size on this screen to 0.3 micron
  870   Fri Aug 22 13:58:39 2008 SharonSummary Trend of the Wiener TF
In order to understand if we really need an adaptive filter, I used old data of MC_L and the accelerometers and seismometer to see if the Wiener (ideal) TF between MC_L and the others really changes all the time.
Two tests I made:
  • Compare the TF after different segments of time, starting from the same point. Meaning, measuring the TF after 5,10,15,20... minutes, looking when and if the TF stablizes (stops changing).
  • Compare the TF between same-length segments, from different times. Meaning, comparing for example 2 segments of 10 minutes taken from different times.

Results:
  • As you can see in the attached PDF, the changes start being minor after 200,000 data points, which correspont to 200,000/256 s, which is approximately 13 minutes.
    If you look at the PDF file, it is arranged from shorter times to longer in the order of: 3, 6, 13, 26 and 39 minutes.

  • As expected, the TF between different segmants of the same length is not completely the same. Again, you can look at the attached PDF.
    Sorry the titles are the same. Each 2 consecutive pages represent the same length of segment in different times. The order of segment's lengths is: 3, 13, 26 and 39 minutes

How do I explain what's going on?

Since the Wiener filter finds the correlation matrix between the data and the noise signals, it will maintain some kind of familiar shape when we don't add a significant amount of unusual data. I am assuming that if I had looked at longer time periods, we could see a more significant change in the TF in time. When looking at different times, the average noise is likely to be different which can explain the change in the correlation matrix and the TF.

To sum up

I think we should give adaptive filtering a go.
Attachment 1: Same_start_time1.pdf
Same_start_time1.pdf Same_start_time1.pdf Same_start_time1.pdf Same_start_time1.pdf Same_start_time1.pdf
Attachment 2: same_length.pdf
same_length.pdf same_length.pdf same_length.pdf same_length.pdf same_length.pdf same_length.pdf same_length.pdf same_length.pdf
  869   Fri Aug 22 10:39:41 2008 JenneUpdateSUSTaking Free Swinging spectra of PRM, SRM, ITMX, BS
I'm taking free swinging spectra of PRM, SRM, ITMX and BS, so I've turned off their watchdogs for now. I should be done around 11:15am, so I'll turn them back on then.
  868   Thu Aug 21 18:13:24 2008 ranaUpdateIOOMC WFS Control signals not responsible for lock losses
This is a 4 hour, second-trend of the MC WFS error and control signals.

There is no sign that the MC loses lock because of feedback signal saturations.
Attachment 1: Untitled.png
Untitled.png
  867   Thu Aug 21 17:55:14 2008 ranaHowToIOOMC WFS DC Offsets
I ran the McWFS_dc_offsets script to trim out the DC offsets on the MC WFS DC signals.

Rob says "who cares?"
  866   Thu Aug 21 16:28:59 2008 steveConfigurationSAFETYsafety glasses required
I just opened the MOPA shutter so PA is warming up.

High power beam path is cleared and laser safety glasses required.
  865   Thu Aug 21 10:24:20 2008 steveConfigurationSAFETYlaser safe mode condition
The MOPA and PSL shutters are closed.
Manual beam blocks are in place.
Enclosure interlock is enabled.
No other high power laser is in operation.

We are in laser safe of operation for visiting students from Japan

NO safety glasses required
  864   Wed Aug 20 18:09:48 2008 YoichiUpdateIOOMC still unlocks
Being suspicious of FSS PC path as the culprit of the MC unlocks, I opened the FSS box and connected a probe to the TP7,
which is a test point in the PC path (before high voltage amplifier).
The signal is routed to an unused fast DAQ channel in the IOO rack. It is named C1:IOO-MC_TMP1 and recorded by the frame
builder. You can use this channel as a generic test DAQ channel later.

By looking at the attachment, the PC path (C1:IOO-MC_TMP1) goes crazy at the same time as other channels. So probably
it is not the trigger for the MC unlock.

Then I noticed the WFS signals drift away just before the unlock as shown in the attached plot. So now the WFS is the
main suspect.
Rob tweaked MC1 pitch to center the WFS QPDs while the MC is not locked. It improved the shape of the MC reflection.
However, the sudden MC unlock still happens. We then lowered the WFS gain from 0.5 to 0.3. Did not change the situation.
It looks like the MC length loop starts oscillating after the WFS signals drift away.
We will measure the WFS and MC OPLTF to see the stability of the loops tomorrow.

Attachment 1: MC-unlock.png
MC-unlock.png
  863   Wed Aug 20 17:02:01 2008 SharonUpdate More FIR to IIR
I tested another method for converting from FIR to IIR other than the 2 mentioned in post 841.
I got this one from Yoichi, called poles fitting, you can read about it more if you want at: http://www.rssd.esa.int/SP/LISAPATHFINDER/docs/Data_Analysis/DA_Six/Heinzel.pdf.

Seems it's not doing much good for us though.

I am attaching a PDF file with the plots, which have N=50,100,600,1000, respectivaly.
Attachment 1: polefit1.pdf
polefit1.pdf polefit1.pdf polefit1.pdf polefit1.pdf
  862   Wed Aug 20 13:23:32 2008 robUpdateLockingDRMI locked

I was able to lock the DRMI this afternoon. All the optical levers have been centered.
  861   Wed Aug 20 12:39:11 2008 EricSummaryCamerasWeekly Summary
I attempted to model the noise produced by the mirror defects in the ETMX images, in order to better assure that the fit to the beam Gaussian in these images is actually accurate. My first attempt involved treating the defects as random Gaussians which were scaled by the power of the beam's Gaussian. This didn't work at all (it didn't really look like the noise on the ETMX), and resulted in very different behavior from the fitting software (it fit to one of the noise peaks, instead of the beam Gaussian). I'll try some other models another time.

I made a copy of the ezcaservo source code and added options to it that allow the addition of minimum value, maximum value, and slew rate limits. This should allow the camera code to servo on ITMX without accidently driving the mirror too far or too fast. In order to get the code to recompile, I had to strip out part of the servo that changed the step value based on the amount of time that had elapsed (it relied on some GDS libraries and header files). Since the amount of time that passes is reasonably constant (about 2-3 steps per second) and the required accuracy for this particular purpose isn't extremely high, I didn't think it would matter very much.

I put together two MATLAB functions that attempt to convert pixel position in an image to actual position in real space. The first function takes four points that have known locations in real space (with respect to some origin which the camera is pointing at) and compares them to where those 4 points fall in the image. From the distortion of the four points, it calculates the three rotational angles of the the camera, as well as a scaling factor that converts pixels to real spatial dimensions. The second function takes these 4 parameters and 'unrotates' the image, yielding the positions of other features in the image (though they must be on the same flat plane) in real space. The purpose of this is to allow the cameras to provide positions in terms of physically meaningful units. It should also decouple the x and y axes so that the two dimensions can be servo'd on independently. Some results are attached; the 'original' image is the image as it came out of the camera (units in pixels), while the 'modified' image is the result of running the two functions in succession. The four points were the corners of the 'restricted access' sign and of the TV screen, while the origin was taken as the center of the sign or the TV. The accuracy of the transformation is reasonably good, but seems to depend considerably on assuring that the origin chosen in real space matches the origin in the image. To make these the same, they will be calculated by taking the intersections of the 2 lines defined by 2 sets diagonal points in each image. The first function will remain in MATLAB, since it only needs to be run once each time the camera is moved. The second function must be ported to C since the transformation must be done in realtime during the servo.

Joe and I attempted another scan of the PMC this morning. We turned the laser power down by a factor of ~50 (reflection off of the unlocked PMC went from ~118 to ~2.2) and blocked one beam in the MZ. We scanned from 40 V to 185 V ( -1 to -4.25 on the PZT ramp channel) with periods of 60 seconds and 10 seconds. In both cases, thermal effects were still clearly visible. We turned the laser power down by another factor of 2 (~1 on the PMC reflection channel), and did a long scan of 300 seconds and a short scan of 10 seconds. The 10 second scan produced what may be clean peaks, although there was clear digitization noise, while the peaks in the 300 second scan showed thermal effects. I've yet to actually analyze the data closely, however.
Attachment 1: OriginalSignImage.png
OriginalSignImage.png
Attachment 2: ModifiedSignImage.png
ModifiedSignImage.png
Attachment 3: OriginalTVImage.png
OriginalTVImage.png
Attachment 4: ModifiedTVImage.png
ModifiedTVImage.png
  860   Wed Aug 20 12:04:47 2008 JenneUpdateSUSBetter diagonalization of PRM input matrix
The values here should replace those in entry #851 from yesterday.

After checking the results of the input matrix diagonalization, I have determined that Sonia's method (described in LIGO-T070168) is more effective at isolating the eigenmodes than Shihori's method (LIGO-T040054).

So, the actual new PRM input matrices are as follows:

POSPITYAW
UL0.96781.0000.7321
UR1.0000.8025 -0.9993
LR0.7235 -1.1230 -1.0129
LL0.6648 -1.04521.0000


Attached are plots of the spectra of the eigenmodes, using both Shihori's and Sonia's methods. Note that there isn't a good way to get the side peak out of the eigenmodes.

I've put these into the SUS-PRM MEDM screen.
Attachment 1: PRM_Eigmodes_shihori.png
PRM_Eigmodes_shihori.png
Attachment 2: PRM_Eigmodes_sonia.png
PRM_Eigmodes_sonia.png
  859   Wed Aug 20 11:50:10 2008 JohnSummaryComputersStripTools on op540m

To restart the striptools on op540m:

cd /cvs/cds/caltech/scripts/general/

./startstrip.csh
  858   Wed Aug 20 11:42:49 2008 JohnSummaryComputerspdftk
I've installed pdftk on all the control room machines.

http://www.pdfhacks.com/pdftk/
  857   Tue Aug 19 19:14:17 2008 YoichiConfigurationDAQFixed C1:IOO-MC_RFAMPDDC
Yoichi, Rob

C1:IOO-MC_RFAMPDDC, which is a PD at the transmission port of the MC, was not recording sensible values.
So I tracked down the problem starting from the centering of the beam on the PD.
The beam was hitting the PD properly. The DC output BNC on the PD provided +1.25V output when the light was
falling on the PD. The PD is fine.
The flat cable from the PD runs to the IOO rack and fed into the LSC PD interface card.
The output from the interface card is connected to a VMIC3113A DAQ card, through cross connects.
The voltages on the cross connects were ok.
The VMIC3113A was controlled by an EPICS machine (c1iool0). So it provides only a slow channel.
By looking at C1IOOF.ini and tpchn_C1.par, I figured that C1:IOO-MC_RFAMPDDC is using chnnum=13639 in the RFM
network and it is named C1:IOO-ICS_CHAN_15 in the .par file. So it is reading values from the ICS DAQ board.
Actually nothing was connected to the channel 15 of the ICS board and that was why C1:IOO-MC_RFAMPDDC was reading
nothing. So I took the PD signal from the cross connect and hooked it up to the Ch15 of the ICS DAQ through
the large black break out box with 4-pin LEMOs. Now C1:IOO-MC_RFAMPDDC reads the DC output of the PD.
I also put an ND filter in front of the RFAMPD to avoid the saturation of the ADC. The attenuation should have been done
electronically, but I was too lazy. Since the ND filter changes the Stochmon values, someone should remove it and reduce the
gain of the LSC PD interface accordingly.
  856   Tue Aug 19 18:55:41 2008 YoichiUpdateIOOMC unlock study update
In entry 849, I reported that the MC transmitted power drops before the sudden unlocks.
However, because C1:IOO-MC_TRANS_SUM is a slow channel, we were not sure if we can believe the timing.
So I wanted to use C1:IOO-MC_RFAMPDDC, which is a fast channel, to monitor the transmitted light power.
However, this channel was broken. So I fixed it. Details of the fixing work is reported in another entry.
The attached plot shows a recent unlock event. It is clear that in the fast channel (i.e. C1:IOO-MC_RFAMPDDC),
there is no delay between the drop of the MC power and the crazy behavior of control signals.
So it was concluded that the apparent precedence of the MC power drop in the slow channels (i.e. C1:IOO-MC_TRANS_SUM)
is just an artifact of timing inaccuracy/offset of the slow epics channels.

Sometime around 5PM, the MC started to be unwilling to even lock. It turned out that the PC drive of the FSS was going
crazy continuously. So I changed the normal values of the common gain and the fast gain, which the mcup script uses.
Now with this new setting, the MC locks happily, but still keeps unlocking.
Attachment 1: MC-unlock.png
MC-unlock.png
  855   Tue Aug 19 17:15:34 2008 SharonUpdate MEDM
I plugged in the gains I got for the accelerometers in the accelerometers' filters in the PEM screen of the adaptive filter
  854   Tue Aug 19 17:00:19 2008 SharonUpdate Wiener TF calibration - update
This is an update for post 814

I added the calibration gains I got for the accelerometers (I realize I am just calibrating the accelerometers to themselves and this is not m/m exactly since we don't really know which accelerometer is doing exactly what we want it to do. However, since we are talking on relative small numbers, this shouldn't really change much).


I also added another missing gain for the seismometer. Rana has previously installed a 4300 ohm resistor in the seismometer, which changed the gain to 4300/(4300+5000) = 0.46 (this is from the manual). Moreover, there is a gain of 100 on the SR560. This comes up to an extra gain of 46, meaning multiplying the seismometer's counts by 1/46.
Attachment 1: m_per_m.png
m_per_m.png
  853   Tue Aug 19 14:25:38 2008 SharonUpdatePEMAccelerometer's calibration - update
Goal - Make sure the accelerometers are calibrated among themselves (have the same power spectrum when they are all together reading the same movements).

What I did - took the accelerometers off their usual X Y Z setting and set the 3 MC2's and 3 MC1's next to each other covered by a box.
Then I brought MC2 X to MC1 X and placed them in a box so I have a referance between the 2 groups.

Result - Seems MC1 accelerometers are much alike and have the same power spectrum when placed together for all frequencies. MC2 accelerometers seem to do the same until approximately 30 Hz. (decided not to correct for that since we don't really care about the accelerometers in such high frequencies).

When comparing the 2 X's, they also seemed to be almost perfectly correlated. I chose the gain by dividing the two and finding the mean of that in the range of 2 to 30 Hz. After correcting for all the accelerometers, I matched the gains of each group to its X accelerometer.

You can see the plots, taking into consideration that the groups were never together (pretty messy getting the cables all around).

Here are the numbers, when the MC2 and MC1 gains are calculated by comparing them to their X direction.

gain MC1 X_over_MC2 X=

1.0870


gain_MC2_Y =

0.9944


gain_MC2_Z =

0.9479


gain_MC1_Y =

1.0394


gain_MC1_Z =

0.9149
Attachment 1: acccalibafter.png
acccalibafter.png
Attachment 2: acccalibbefore.png
acccalibbefore.png
  852   Tue Aug 19 13:34:58 2008 josephbConfigurationComputersSwitched c1pem1, c0daqawg, c0daqctrl over to new switches
Moved the Ethernet connections for c1pem1, c0daqawg, and c0daqctrl over to the Netgear Prosafe switch in 1Y6, using new cat6 cables.
  851   Tue Aug 19 13:12:55 2008 JenneUpdateSUSDiagonalized PRM Input Matrix
NOTE: Use the values in elog #860 instead (20Aug2008)

Using the method described in LIGO-T040054-03-R (Shihori's "Diagonalization of the Input Matrix of the Suspension System"), I have diagonalized the input matrices for the PRM.

Notes about the method in the document:
  • Must define the peak-to-peak voltage (measured via DataViewer) to be NEGATIVE for PitLR, PitLL, YawUR, YawLR, and POSITIVE for all others
  • As Osamu noted in his 3 Aug 2005 elog entry, all of the negative signs in equations 4-9 should all be plus.

New PRM Input Matrices:

POSPITYAW
UL1.0001.0001.000
UR1.18771.0075 -1.0135
LR0.8439 -0.9425 -0.9653
LL0.9684 -1.05001.0216
  850   Tue Aug 19 10:36:34 2008 SharonUpdate Calibrating accelerometers
I took apart the accelerometers near MC1 and MC2.
The 2 sets of 3 accelerometers are now covered by a box on the floor. Please try not to move them... I will place it all back once I am done calibrating.
  849   Mon Aug 18 22:47:12 2008 YoichiUpdateIOOMC unlock study
As rob noted, the MC keeps unlocking in a few minutes period.
I plotted time series of several signals before unlocks.
It looks like the MC alignment goes wrong a few hundred msec before the unlock (the attached plot is only one example, but all unlocks
I've looked so far show the same behavior).
I will look for the cause of this tomorrow.

The horizontal axis of the plot is sec. The data values are scaled and offset-removed appropriately so that all curves are shown
in a single plot. Therefore, the vertical axis is in arbitrary units.
Attachment 1: MC-Unlock.png
MC-Unlock.png
ELOG V3.1.3-